OpenStudy (anonymous):

1/(3x+1)^2 from 1 to inf

OpenStudy (anonymous):

are you looking for \[\int_1^{\infty}\frac{dx}{(3x+1)^2}\]?

OpenStudy (anonymous):

anti - derivative is easy enough. use a u-sub \[u=3x+1,\frac{1}{3}du=dx\] to get \[\frac{1}{3}\int \frac{1}{u^2}du=-\frac{1}{3}\frac{1}{u}=-\frac{1}{9x+3}\]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

now evaluate at r and 1, you get \[-\frac{1}{9r+3}+\frac{1}{12}\] and then take \[\lim_{r \rightarrow \infty}-\frac{1}{9r+3}+\frac{1}{12}=0+\frac{1}{12}=\frac{1}{12}\]